OpenStudy (anonymous):

Two jets are beginning their flight at the Louisville International Airport. One starts at 1:00 p.m. and flies east at 395 mph. Another starts one hour later and flies west at 400 mph. In how many hours after the first plane left, will they be 2,000 miles apart? Round to the nearest tenth of an hour.

OpenStudy (anonymous):

just did this one for school :) 2, Two jets are beginning their flight at the Louisville International Airport. One starts at 1:00 p.m. and flies east at 395 mph. Another starts one hour later and flies west at 400 mph. In how many hours after the first plane left, will they be 2,000 miles apart? Round to the nearest tenth of an hour. After 1 hour they are 395 miles apart For every hour after that they will = (395+400) = 795 miles apart. Accounting for the first 395 miles, you need to know how long it will take to get (2000-395) = 1605 miles apart. Divide 1605 by 795 and get 2.018 hours. Remember to add 1 hour for the part where only 1 plane was in the air, and the total is 3.018 to the nearst 1/10 hour this is 3. Answer is B